# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

### v3

@@ 8 Acknowledgments
A. Bagheri acknowledges the support from the Swiss National Science Foundation by grant number P500PT_214435. A. Berne is supported by the Future Investigators in NASA Earth and Space Science and Technology (FINESST) Program (80NSSC22K1318). We also acknowledge the two anonymous reviewers for their constructive suggestions which helped improvement of the paper.
@@ Appendix A Viscoelastic models
Figure 13: Difference between different tidal observables as due to the use of the Maxwell and Andrade rheology: (a) tidal heating ($\Delta\dot{E}$), (b) maximum radial displacements ($\Delta R$), $k_{2}$ (c) and $h_{2}$ (d) Love numbers ($\Delta k_{2}$ and $\Delta h_{2}$) and associated phase lags ($\Delta\epsilon_{k_{2}}$ and $\Delta\epsilon_{h_{2}}$) (e, f). The choice of the viscoelastic model between Andrade and Maxwell’s models is typically more important as the tidal forcing periods become closer to the intrinsic relaxation time of the body. The choice of viscoelastic models for Enceladus can be potentially important because the orbital frequency of Enceladus (32.9 hr) is close to the estimated relaxation time for ice ($\tau_{M}=\eta_{ice}/\mu_{ice}$). The four narrow regions constrained by the red lines in panels (a) and (b) indicate the estimated tidal heat and the associated core viscosities (see Figure 1).
@@
Figure 14: Effect of changing the Andrade parameter $\alpha$ on the tidal response of Enceladus. Right and left panels show the tidal response parameters for models of tidal dissipation dominantly in the shell and in the core, respectively.
Appendix B Sensitivity of libration to elastic response
@@
As highlighted in Section 4, if the ice shell is sufficiently thick as inferred for Enceladus, the libration amplitude is mostly determined by the moment of inertia and therefore, the thickness and density of the shell. We further explore the variations of the shell libration amplitude and the period of free libration as functions of the product of the shell density and shell thickness as a representative of its moment of inertia (Figure 15). We also show the same quantities with respect to the ratio of shell thickness to the shell density for better understanding of the behavior. As shown in the figure, the libration amplitude has dramatically large values for a certain value of $\rho_{shell}\times D_{shell}$ where the natural frequency of the shell equals the orbital period of Enceladus.
@@
Figure 15: (a) Natural libration period (hr) and (b) the amplitude of forced physical libration of Enceladus as functions of the product of shell thickness and density as a representative of its shell MoI. (c,d) Same quantities as a function of the ratio of shell thickness and shell density. The plots are produced for fine grids of shell thickness and densities in ranges [1-60] km and [400-900] kg/m3. Colorbar indicates the associated shell thickness.
@@
Figure 17: Variation of the amplitude of the libration as a function of ice shell thickness and the effective viscosities of shell (a) and the core (b). In panel (a), the core viscosity is assumed equal to $\rm 10^{20}~Pa.s$. The region between the red lines shows the libration amplitude spanning the same range in Figure 3. Other interior structure parameters are the same as in Figure 2. At $\rm D_{shell}\approx 5$ km, the natural free libration period of the shell approaches the orbital period of Enceladus (for $\rho_{shell}$ fixed here to 900 kg/m3.). The shell thickness and amplitude of libration are plotted in logarithmic scales.
Appendix C Derivation of tidal heating partitioning between the shell and the core
We can compute the tidal heating in the specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body as a result of tidal perturbations. Note that Equation (5) cannot be used for the core or the surrounding layers simply by replacing $k_{2}/Q$ or $\Im(k_{2})$ of the whole body by the tidal Love numbers of the core or the layers computed independently of the rest of the body. Here, we use equations (33) and (37) in Tobie et al. (2005) to derive the contribution of the core to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (C1)
$\dot{E}_{core}=-\frac{42\pi}{5}\omega^{5}e^{2}R_{E}^{4}\int_{0}^{R_{core}}H_{\mu}\Im{(\mu(r))dr}.$ (C2)
$\int_{0}^{R_{core}}H_{\mu}\Im{(\mu(r))dr}=\frac{5R_{core}}{4\pi G}\Im{(\hat{k}_{2}^{core}}).$ (C3)
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (C4)
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (C5)
$\hskip 26.46085pt-\frac{21}{2G}\omega^{5}e^{2}R_{E}^{4}\bigg{(}R_{E}\Im{(k_{2})}-R_{core}\Im{(\hat{k_{2}}^{core})}\bigg{)}.$ (C6)
Here, $R_{E}$ and $R_{OB}$ refer to the radius at Enceladus’s surface at the bottom of the ocean. Equations (C2), (C5), and (C) provide a method to calculate the partitioning of global tidal heating for a given interior structure model. These equations require a priori information, assumptions, or independent constraints on the rheological and structural properties of the core to compute the shear dissipation in the core.
This method can be used in a forward model by assumptions on the core properties or in the context of an inverse problem based on a sampling approach. However, we consider requirements on the geophysical measurement that can not directly depend on the interior model parameters. Thus, we extend our analysis to derive the tidal heating on the basis of potentially observable quantities. Quantifying the partitioning of the heat in the icy body is equal to calculating the contribution of the core and the shell to the imaginary part of the gravitational Love number $\Im{(k_{2})}$, as indicated from Equation (5). Beuthe (2019) uses membrane mechanics to obtain the contribution of each layer within the body to the imaginary part of the tidal Love number:
$\Im{(k_{2})}=\left|\frac{k_{2}+1}{k_{2}^{o}+1}\right|^{2}\Im{(k_{2}^{o})}-\xi_{2}|h_{2}|^{2}\Im{(\Lambda_{T})},$ (C7)
$\dot{E}_{\text{core}}=-\frac{5}{2}\frac{\omega R}{G}\left|\frac{k_{2}+1}{k_{2}^{0}+1}\right|^{2}\Im(k_{2}^{0})\langle|U_{2}^{(T)}|^{2}\rangle,$ (C8)
$\dot{E}_{\text{shell}}=\frac{5}{2}\frac{\omega R}{G}\xi_{2}\,|h_{2}|^{2}\Im(\Lambda_{T})\,\langle|U_{2}^{(T)}|^{2}\rangle,$ (C9)
where,
$\langle|U_{2}^{(T)}|^{2}\rangle=(\frac{21}{5})(\omega R)^{4}e^{2}.$ (C10)
The fluid crust gravitational Love number can be computed from Equation D.2 in Appendix D and Table 9 in Beuthe (2019). Once the tidal heating rate in the shell is computed, the heat produced in the core can be computed by subtracting the heating in the shell from the total heat. Following Beuthe (2015), the term $\Lambda_{T}$ in Equation (C8) is the complex spring constant of the shell and is computed as:
$\Lambda_{T}=\Lambda_{M}+\Lambda_{\rho}+\Lambda_{\chi}+\Lambda_{\omega}$ (C11)
$\Lambda=f_{\mu}\hat{\mu}\epsilon,$ (C12)
$\Lambda_{\rho}=f_{\rho}\frac{\delta\rho}{\rho}\epsilon.$ (C13)
$f_{\mu}=\frac{8(1+{\nu})}{(5+{\nu})},$ (C14)
$f_{\rho}=\frac{1+5{\nu}}{5+{\nu}},$ (C15)
and,
$\hat{\mu}=\frac{\mu}{\rho gR},$ (C16)
$\nu=\frac{3K-2\mu}{6K+2\mu}.$ (C17)
This formulation provides a means for obtaining the tidal heating in the shell by measuring the tidal Love number $h_{2}$ (real and imaginary parts) and independent constraint on the imaginary part of the complex shear modulus of the shell, $\Lambda_{T}$, as shown in Equation (C8). Here, we extend the formulation to develop expressions for the partitioning of heat between the shell and the core, which do not require a priori knowledge of the shell rheology. Equation 125 in Beuthe (2015) provides an expression between the displacement and potential Love numbers, the rigidity of the shell, and the density structure of the body:
$k_{2}+1=(1+\Lambda_{T})h_{2}-5\frac{\delta\rho}{\rho}\epsilon.$ (C18)
$\Im{(\Lambda_{T})}=\Im\bigg{(}\frac{1}{h_{2}}{(k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon)}\bigg{)},$ (C19)
Combining Equations (C8) and (C19), we can rewrite the expression of tidal heating in the shell as:
$\dot{E}_{\text{shell}}=\frac{105}{10}\frac{(\omega R)^{5}}{G}\xi_{2}\,|h_{2}|^{2}\bigg{[}\Im\bigg{(}{\frac{1}{h_{2}}\big{(}k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon\big{)}}\bigg{)}\bigg{]}\,e^{2}.$ (C20)
This formulation provides the tidal heating in the shell by measurements of the real and imaginary parts of the tidal Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in Equation (C8). Tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
@@ Appendix E Probability Density Function (PDF) of interior model parameters and correlation
Figure 18: Exploration of the key geophysical interior model parameters corresponding to with Figure 4. Multiple peaks observed in the PDFs of the viscosities of ice shell and the core correspond to separate tidal dissipation regimes that can match the inferred global heat loss. These peaks, and the correlation observed between the viscosities of the shell and rock correspond to the regions indicated in Figure 1. Shear moduli are presented in GPa, $\rm D_{shell}$ and $\rm R_{core}$ are in km, and other parameters are presented in their SI units. Ice shell thickness, and shear moduli are plotted in logarithmic axes.
@@
Assuming the same set of measurements, but limiting the viscosity ranges of the shell and core to $10^{11}<\eta_{shell}<10^{15}$  Pa.s and $10^{11}<\eta_{core}<10^{20}$ Pa.s, and separating the two sets of models based on the two dissipation scenarios, we obtain the interior parameters shown in Figures 17a and 17b.
We compare the amplitude and phase lags of $h_{2}$ and $l_{2}$ in Figure 19 for the two tidal dissipation scenarios. The 2-$\sigma$ standard deviations are indicated in this figure.
Figure 19: Vertical and horizontal displacement Love numbers ($|h_{2}|$, $|l_{2}|$) and phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) are presented for the core-dominated and shell-dominated tidal dissipation. The heat flux is assumed to be $\rm 15~GW<\dot{E}<40~GW$, covering estimates in Hemingway et al. (2018) and Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). The top axis in the plots for $h_{2}$ and $l_{2}$ indicate the associated maximum radial and maximum easterly displacements corresponding to the values of the Love number in the bottom axis, respectively.
When considering the constraints on the geodetic observables of Park et al. (2024) and separating the interior parameters for the two tidal dissipation scenarios, we obtain constraints on the interior parameters and the correlations between them, as shown in Figures 18b and 18c.
@@
Figure 17b: Exploration of the correlation between the model parameters associated with Figure 5 for the case tidal dissipation is dominated in the shell. Units and plotting conventions are similar to Figure 18.
Figure 18a: Exploration of the correlation between the model parameters associated with Figure 7. The ranges of the plausible interior parameters are shown in Table 4. Blue and red colors indicate the lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
Figure 18b: Model parameters for the case where at least 90% of the tidal heating is focused in ice shell associated with Figures 7 and 18a. The tidal heating in the shell is calculated using the formulation by (Beuthe, 2019) (Equation (C8)). Black and yellow colors indicate lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
Figure 18c: Model parameters for the case where at least 90% of the tidal heating is focused in the rocky core associated with Figures 7 and 18a. The tidal heating in the core is calculated by subtracting the heat in the shell from the total tidal heating computed based on Equation (5). Blue and red colors indicate the lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.

### v5

@@ 7 Acknowledgments
A. Bagheri acknowledges the support from the Swiss National Science Foundation by grant number P500PT_214435. A. Berne is supported by the Future Investigators in NASA Earth and Space Science and Technology (FINESST) Program (80NSSC22K1318). We acknowledges insightful discussions with Francis Nimmo, Michael Efroimsky, Mikael Beuthe and Sander Goossens. We also acknowledge the two anonymous reviewers for their constructive suggestions which helped improvement of the paper.
@@ Appendix A Viscoelastic models
Figure 13: Difference between different tidal observables as due to the use of the Maxwell and Andrade rheology: (a) tidal heating ($\Delta\dot{E}$), (b) maximum radial displacements ($\Delta R$), $k_{2}$ (c) and $h_{2}$ (d) Love numbers ($\Delta k_{2}$ and $\Delta h_{2}$) and associated phase lags ($\Delta\epsilon_{k_{2}}$ and $\Delta\epsilon_{h_{2}}$) (e, f). The four narrow regions constrained by the red lines in panels (a) and (b) indicate the estimated tidal heat (see Figure 2).
@@
Figure 14: Effect of changing the Andrade parameter $\alpha$ on the tidal response of Enceladus. Left and Right panels show the tidal response parameters for models of tidal dissipation dominantly in the shell and in the core, respectively.
Appendix B Derivation of partitioning of tidal heating between the shell and the core
We can compute the tidal heating in the specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body as a result of tidal perturbations. Note that equation (5) cannot be used for the core or the surrounding layers simply by replacing $k_{2}/Q$ or $\Im(k_{2})$ of the body by the tidal Love numbers of the core or the layers computed independently of the rest of the body. Here, we use equations (33) and (37) in Tobie et al. (2005) to derive the contribution of the core to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (B1)
$\dot{E}_{core}=-\frac{42\pi}{5}\omega^{5}e^{2}R_{E}^{4}\int_{0}^{R_{core}}H_{\mu}\Im{(\mu(r))dr}.$ (B2)
$\int_{0}^{R_{core}}H_{\mu}\Im{(\mu(r))dr}=\frac{5R_{core}}{4\pi G}\Im{(\hat{k}_{2}^{core}}).$ (B3)
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (B4)
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (B5)
$\hskip 26.46085pt-\frac{21}{2G}\omega^{5}e^{2}R_{E}^{4}\bigg(R_{E}\Im{(k_{2})}-R_{core}\Im{(\hat{k_{2}}^{core})}\bigg).$ (B6)
Here, $R_{E}$ and $R_{OB}$ refer to the radius at Enceladus’s surface at the bottom of the ocean. equations (B2), (B5), and (B) provide a method to calculate the partitioning of global tidal heating for a given interior structure model. These equations require a priori information, assumptions, or independent constraints on the rheological and structural properties of the core to compute the shear dissipation in the core.
This method can be used in a forward model by assumptions on the core properties or in the context of an inverse problem based on a sampling approach. However, we consider requirements on the geophysical measurement that can not directly depend on the interior model parameters. Thus, we extend our analysis to derive the tidal heating on the basis of potentially observable quantities. Quantifying the partitioning of the heat in the icy body is equal to calculating the contribution of the core and the shell to the imaginary part of the gravitational Love number $\Im{(k_{2})}$, as indicated from equation (5). Beuthe (2019) uses membrane mechanics to obtain the contribution of each of the interior layers to the imaginary part of the tidal Love number:
$\Im{(k_{2})}=\left|\frac{k_{2}+1}{k_{2}^{o}+1}\right|^{2}\Im{(k_{2}^{o})}-\xi_{2}|h_{2}|^{2}\Im{(\Lambda_{T})},$ (B7)
$\dot{E}_{\text{core}}=-\frac{5}{2}\frac{\omega R}{G}\left|\frac{k_{2}+1}{k_{2}^{0}+1}\right|^{2}\Im(k_{2}^{0})\langle|U_{2}^{(T)}|^{2}\rangle,$ (B8)
$\dot{E}_{\text{shell}}=\frac{5}{2}\frac{\omega R}{G}\xi_{2}\,|h_{2}|^{2}\Im(\Lambda_{T})\,\langle|U_{2}^{(T)}|^{2}\rangle,$ (B9)
where,
$\langle|U_{2}^{(T)}|^{2}\rangle=(\frac{21}{5})(\omega R)^{4}e^{2}.$ (B10)
The fluid crust gravitational Love number can be computed from equation D.2 in Appendix D and Table 9 in Beuthe (2019). Once the tidal heating rate in the shell is computed, the heat produced in the core can be computed by subtracting the heating in the shell from the total heat. Following Beuthe (2015), the term $\Lambda_{T}$ in equation (B8) is the complex spring constant of the shell and is computed as:
$\Lambda_{T}=\Lambda_{M}+\Lambda_{\rho}+\Lambda_{\chi}+\Lambda_{\omega}.$ (B11)
$\Lambda=f_{\mu}\hat{\mu}\epsilon,$ (B12)
$\Lambda_{\rho}=f_{\rho}\frac{\delta\rho}{\rho}\epsilon.$ (B13)
$f_{\mu}=\frac{8(1+{\nu})}{(5+{\nu})},$ (B14)
$f_{\rho}=\frac{1+5{\nu}}{5+{\nu}},$ (B15)
and,
$\hat{\mu}=\frac{\mu}{\rho gR},$ (B16)
$\nu=\frac{3K-2\mu}{6K+2\mu}.$ (B17)
This formulation provides a means for obtaining the tidal heating in the shell by measuring the tidal Love number $h_{2}$ (real and imaginary parts) and independent constraint on the imaginary part of the complex shear modulus of the shell, $\Lambda_{T}$, as shown in equation (B8). Here, we extend the formulation to develop expressions for the partitioning of heat between the shell and the core, which do not require a priori knowledge of the shell rheology. equation 125 in Beuthe (2015) provides an expression between the displacement and potential Love numbers, the rigidity of the shell, and the density structure of the body:
$k_{2}+1=(1+\Lambda_{T})h_{2}-5\frac{\delta\rho}{\rho}\epsilon.$ (B18)
$\Im{(\Lambda_{T})}=\Im\bigg(\frac{1}{h_{2}}{(k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon)}\bigg),$ (B19)
Combining equations (B8) and (B19), we can rewrite the expression of tidal heating in the shell as:
$\dot{E}_{\text{shell}}=\frac{105}{10}\frac{(\omega R)^{5}}{G}\xi_{2}\,|h_{2}|^{2}\bigg[\Im\bigg({\frac{1}{h_{2}}\big(k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon\big)}\bigg)\bigg]\,e^{2}.$ (B20)
This formulation provides the tidal heating in the shell by measurements of the real and imaginary parts of the tidal Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in equation (B8). Tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
Appendix C Sensitivity of libration to interior properties
@@
As highlighted in Section 3, if the ice shell is sufficiently thick as inferred for Enceladus, the libration amplitude is mostly determined by the moment of inertia and therefore, the thickness and density of the shell. We further explore the variations of the shell libration amplitude and the period of free libration as functions of the product of the shell density and shell thickness as a representative of its moment of inertia (Figure 15). As shown in the figure, the libration amplitude has dramatically large values for a certain value of $\rho_{shell}\times D_{shell}$ where the natural frequency of the shell equals the orbital period of Enceladus.
@@
Figure 15: (a) Natural libration period and (b) the amplitude of forced physical libration of Enceladus as functions of the product of shell thickness and density as a representative of its shell MoI. (c,d) Same quantities as a function of the ratio of shell thickness and shell density. The plots are produced for fine grids of shell thickness and densities in ranges [1-60] km and [400-900] kg/m3. Colorbar indicates the associated shell thickness.
@@
Figure 17: Variation of the amplitude of the libration as a function of ice shell thickness and the effective viscosities of shell (a) and the core (b). In panel (a), the core viscosity is assumed equal to $\rm 10^{20}~Pa.s$. The region between the red lines shows the libration amplitude spanning the same range in Figure 1. Other interior structure parameters are the same as in Figure 3. The shell thickness and amplitude of libration are plotted in logarithmic scales.
@@ Appendix E Probability Density Function (PDF) of interior model parameters and correlation
Figure 18: Exploration of the key geophysical interior model parameters corresponding to with Figure 4. Multiple peaks observed in the PDFs of the viscosities of ice shell and the core correspond to separate tidal dissipation regimes that can match the inferred global heat loss. Shear moduli are presented in GPa, $\rm D_{shell}$ and $\rm R_{core}$ are in km, and other parameters are presented in their SI units. Ice shell thickness, and shear moduli are plotted in logarithmic axes. Black and yellow colors indicate the lowest and highest likelihoods.
@@
Assuming the same set of measurements, but limiting the viscosity ranges of the shell and core to $10^{11}<\eta_{shell}<10^{15}$  Pa.s and $10^{11}<\eta_{core}<10^{20}$ Pa.s, and separating the two sets of models based on the two dissipation scenarios, we obtain the interior parameters shown in Figures 19a and 19b. We compare the amplitude and phase lags of $h_{2}$ and $l_{2}$ in Figure 20 for the two tidal dissipation scenarios. The 2-$\sigma$ standard deviations are indicated in this figure.
When using the inferred geodetic quantities by Park et al. (2024) and separating the interior parameters for the two tidal dissipation scenarios, we obtain constraints on the interior parameters and the correlations between them, as shown in Figures 21b and 21c.
@@
Figure 19b: Exploration of the correlation between the model parameters associated with Figure 5 for the case tidal dissipation is dominated in the shell. Units and plotting conventions are similar to Figure 18. Blue and red colors indicate the lowest and highest likelihoods.
Figure 20: Vertical and horizontal displacement Love numbers ($|h_{2}|$, $|l_{2}|$) and phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) are presented for the core-dominated and shell-dominated tidal dissipation. The heat flux is assumed to be $\rm 15~GW<\dot{E}<40~GW$, covering estimates in Hemingway et al. (2018) and Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). The top axis in the plots for $h_{2}$ and $l_{2}$ indicate the associated maximum radial and maximum easterly displacements in the bottom axis, respectively.
Figure 21a: Exploration of the correlation between the model parameters associated with Figure 7. The ranges of the plausible interior parameters are shown in Table 5. Black and yellow colors indicate the lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
Figure 21b: Model parameters for the case where at least 90% of the tidal heating is focused in ice shell associated with Figures 7 and 21a. The tidal heating in the shell is calculated using the formulation by (Beuthe, 2019) (equation (B8)). Darkest and brightest colors indicate lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
Figure 21c: Model parameters for the case where at least 90% of the tidal heating is focused in the rocky core associated with Figures 7 and 21a. The tidal heating in the core is calculated by subtracting the heat in the shell from the total tidal heating computed based on equation (5). Units and plotting conventions are similar to Figure 18.